### v1

# DOEPatch: Dynamically Optimized Ensemble Model for Adversarial Patches Generation

Abstract
Object detection is a fundamental task in various applications ranging from autonomous driving to intelligent security systems. However, recognition of a person can be hindered when their clothing is decorated with carefully designed graffiti patterns, leading to the failure of object detection. To achieve greater attack potential against unknown black-box models, adversarial patches capable of affecting the outputs of multiple-object detection models are required. While ensemble models have proven effective, current research in the field of object detection typically focuses on the simple fusion of the outputs of all models, with limited attention being given to developing general adversarial patches that can function effectively in the physical world. In this paper, we introduce the concept of energy and treat the adversarial patches generation process as an optimization of the adversarial patches to minimize the total energy of the “person” category. Additionally, by adopting adversarial training, we construct a dynamically optimized ensemble model. During training, the weight parameters of the attacked target models are adjusted to find the balance point at which the generated adversarial patches can effectively attack all target models. We carried out six sets of comparative experiments and tested our algorithm on five mainstream object detection models. The adversarial patches generated by our algorithm can reduce the recognition accuracy of YOLOv2 and YOLOv3 to 13.19% and 29.20%, respectively. In addition, we conducted experiments to test the effectiveness of T-shirts covered with our adversarial patches in the physical world and could achieve that people are not recognized by the object detection model. Finally, leveraging the Grad-CAM tool, we explored the attack mechanism of adversarial patches from an energetic perspective.
Index Terms:
Object detection model, adversarial patch, ensemble model.
I Introduction
The rapid development of artificial intelligence has raised major security concerns in several fields, including image processing[1], reinforcement learning[2, 3], natural language processing [4], and unmanned aerial systems[5, 6]. Object detection models in deep learning, in particular, are vulnerable to adversarial samples, which can introduce carefully crafted perturbation[7, 8, 9], patch[10, 11], or camouflage[12, 13] into digital images and cause wrong predictions. This has emerged as a significant threat to the security of object detection models. Recently, a large amount of research has shown that adversarial examples are transferable[14, 15, 16], meaning that adversarial examples designed for one model can mislead other black-box models. Specifically, Cai et al.[17] proposed an adversarial attack method against dense prediction models. The adversarial examples generated by this method can successfully attack multiple object detection and segmentation models. Simultaneously, with the emergence of adversarial patches, recent research has proposed various attack methods based on these patches, which can be directly printed on cardboard[18] and clothing[19], thereby impacting the recognition of object detection models in the physical world. This poses significant threats to areas such as pedestrian monitoring and autonomous driving.
In practical application scenarios, adversarial patches often confront unknown object detection models. However, the existing studies on adversarial patches have mainly emphasized the deployment of patches in the physical world[20] and attacks under multiple viewpoints[21], the problem of the transferability of adversarial patches has been largely ignored. We generated several sets of adversarial samples randomly from the testing dataset. Apparently, the adversarial patches designed specifically for the YOLOv2[22] object detection model, such as AdvPatch[18] and TCEGA[21], are difficult to affect the detection results of other models such as YOLOv3[23], YOLOv4[24], and Faster R-CNN[25](Fig.1). And our generated Dynamically Optimized Ensemble Patch(DOEPatch) was capable of incapacitating all the object detection models.
The various network architectures used in different object detection models exhibit diversity, and the feature maps obtained for the same sample are also distinct. Adversarial patches mainly disrupt the detection of the original target by the object detection model through special feature distributions generated during training. Therefore, adversarial patches tailored to a single object detection model are difficult to interfere with the detection of other unknown models. Although the transferability issue of adversarial samples has been widely studied, and ensemble-based methods have been widely used with significant effects in addressing the transferability problem. However, earlier ensemble models[26, 27] simply combined the outputs of different models without specifically balancing the weight parameters between victim models, which easily led to the optimization of adversarial patches in a singular direction, limiting the attack performance of ensemble models.
Figure 1: The detection results of mainstream object detection models, on the adversarial patches, namely AdvPatch, TCEGA, and DOEPatch. The red bounding box denotes mistaken detection or the failure of the object detector to recognize the human, whereas the green bounding box signifies correct detection.
To improve the transferability of adversarial patches using ensemble models and explore the attack mechanisms for generating such patches, we conduct an energy-based analysis to gain insights into the generation process. We then develop a universal adversarial patches attack framework by combining this analysis with a dynamic optimization approach for an ensemble model. Specifically, the energy-based model[28, 29] generates low-energy samples from random noise inputs and subsequently enhances the quality of generated samples by updating the network parameters. We introduce the concept of energy-based models to reconstruct the energy function for generating adversarial patches targeted toward object detection models. This is achieved by reducing the energy of the “person” category to effectively launch an attack on the object detectors. During the training of adversarial patches, the confidence scores of bounding boxes related to the “person” category are incorporated into the energy-based function. Through iterative optimization, the energy of the generated adversarial patches is gradually increased, while the energy associated with the target category is reduced.
To enhance the generalization of adversarial patches and strengthen their transferability across different object detection models, we utilize the concept of Min-Max adversarial training to construct a dynamically optimized ensemble model. Therefore, the process of adversarial patch generation can be divided into two main stages. The first stage entails internal maximization, wherein the weights of the target object detection model are dynamically adjusted to maximize the overall energy output. The second stage involves external minimization, wherein the adversarial patch is updated to minimize the overall energy output. Finally, by continuously undergoing Min-Max adversarial training and dynamically adjusting the weight parameters of the models, an adversarial patch that is effective against all target object detection models can be generated and achieved a balanced outcome.
In this study, we developed six types of adversarial patches tailored to mainstream one-stage and two-stage object detection models. We conducted a thorough evaluation of the attack effectiveness and transferability of these patches across five object detection models. Our experimental results demonstrate the transferability of the proposed adversarial patches. In addition, we have achieved the ability to influence the recognition of a person by object detection models in the physical world by printing adversarial patches on clothing. Finally, we performed an explanatory analysis of the effects of our adversarial patches on object detection predictions by examining the variations in target category energy and visualizing feature distributions using Grad-CAM[30] on adversarial samples. The main contributions of this work can be summarized as follows:
  • We proposed novel universal adversarial patches, based on an energy function and a dynamically optimized ensemble model, that can be applied to multiple object detection models.
  • The effectiveness and transferability of the proposed adversarial patches were successfully demonstrated on various state-of-the-art object detection models, including physical-world attacks by printing the patches on clothing.
  • We conducted an explanatory analysis of the adversarial patches’ attacks from an energy perspective and visualized feature distributions using Grad-CAM.
The sections are organized as followed: Section II summarizes the related work on ensemble models and adversarial patches. Section III describes the proposed algorithm. Section IV presents the experimental setup, corresponding results, and analysis. Section V provides the conclusion for this paper.
II Related Works
This paper mainly studies the adversarial patch attack method based on the ensemble model, aiming to generate a universal adversarial patch that can attack multiple object detection models simultaneously. Therefore, we mainly introduce the work about ensemble-based attacks and adversarial patch attacks.
II-A Ensemble-based Attacks
Early adversarial attack research primarily focused on digital world attacks against image classification models[31, 32, 33]. By adding imperceptible perturbations to the original images, classification models could be deceived and make incorrect predictions. The resulting images with added perturbation were known as adversarial examples. With the widespread application of adversarial examples, they could be transferred between different architectures, severely affecting the applicability of neural networks. Some efforts have been made to enhance the transferability of adversarial examples by training substitute models[34], but it was challenging to apply this approach to large models and large datasets. Liu et al.[35] pioneered the research on the transferability of large-scale models and datasets by introducing an approach based on the ensemble of multiple target models to generate transferable adversarial examples. Based on the ensembled model attack, by training an adversarial sample that can deceive multiple attacked models, the success rate of attacks against unknown black-box models will also increase. Zheng et al.[36] introduced the Meta Gradient Adversarial Attack (MGAA) framework, which can be seamlessly ensembled with existing gradient-based attack methods to improve the transferability of adversarial examples across different models. Despite the extensive research on adversarial attacks in the field of image classification, there has been a lack of attention on ensemble-based attacks for dense prediction models such as object detection[26, 27]. Meanwhile, most of the existing ensemble model works[37, 38, 39] simply fuse the outputs of all models, neglecting the differences in model structures, which leads to adversarial optimization mostly heading in a single direction, making it difficult for adversarial examples to converge into a form that is effective for all target models. Recently, Cai et al.[17] proposed an attack algorithm based on transfer and query combination, demonstrating that adjusting the weights of the ensemble according to the victim model can further improve the attack performance. The universal adversarial perturbations they obtained can successfully attack various object detection models, such as YOLOv3 and Faster R-CNN, but can only be applied to digital images. Our research focused on the application of ensemble models in adversarial patches, adjusting weight coefficients based on energy loss feedback from different victim models to generate effective patches for all models.
II-B Adversarial Patch Attacks
The adversarial patch was a special form of adversarial samples and had become a popular research object in the field of adversarial attacks. The adversarial patch covered the local area of the image like a sticker pattern, and the victim models could be successfully deceived simply by placing the patch on the object to be tested[10]. How to generate adversarial patches with high attack and migration capabilities was the most important problem at this stage. Wu et al.[19] analyzed the attack performance and transferability of adversarial patches across various datasets, target models, and attack types. Hao et al.[27] utilized critical pixel locations to add adversarial patches, balance the gradients of both detectors during training, and successfully conducted simultaneous attacks on YOLOv4 and Faster R-CNN object detectors. As object detection continues to rapidly develop, there is a growing body of research that focuses on detecting individuals in the physical world. Earlier studies produced adversarial perturbations in digital environments, which were then used to perform physical attacks on actual objects through the creation of adversarial disguises or by directly printing perturbation. Thys et al.[18] had devised adversarial patches that are specifically tailored for object detection systems. These patches, when printed on a cardboard surface, resulted in the first-ever implementation of “stealth” capabilities for humans under physical world object detection models. Hu et al.[21] had developed a novel approach to generate scalable adversarial textures that are suitable for attacking object detection models from multiple angles. Their methods involved utilizing a generator and toroidal cropping strategy to create textures that can be applied to surfaces of various shapes. In contrast to the aforementioned research, our study is centered on the development of universal adversarial patches and the exploration of their transferability and interpretability.
III Methodology
To generate adversarial patches that exhibit effective attacks against mainstream object detectors and demonstrate good transferability, while retaining attack effectiveness in the physical world through printing on fabrics. We investigate the mechanism of generating adversarial patches from the perspective of an energy function. The adversarial patches are first processed by the transformation function and then overlaid on clean samples. The obtained adversarial samples are used as the input of the dynamically optimized ensemble model, and the energy loss output by the ensemble model is ${L_{patch}}$. To enhance the effectiveness of adversarial patches, we introduce two additional components: pixel value constraints and smoothness adjustment. The output of Non-Printability Score[40] loss ${L_{NPS}}$ and smoothness loss ${L_{smooth}}$ are used as another part of the energy function. By integrating the aforementioned three losses, we derive the overall energy loss and subsequently optimize the adversarial patches via gradient descent to minimize this loss. This approach enables the transferability of the adversarial patch across various object detection models. The overall framework is illustrated in the following Fig.2.
Figure 2: Overview of DOEPatch generation framework. The solid line indicates the direction of training the adversarial patch, and the dashed line indicates the direction of backpropagation.
III-A Problem Definitions
Given clean samples ${X}$ with labels ${Y}$, the resulting adversarial examples(${{X_{adv}}{\rm{=}}X+\delta}$), obtained by adding adversarial perturbation ${\delta}$ onto the images, have the potential to cause the object detection model ${f(\cdot)}$ to make an erroneous prediction. To limit the adversarial perturbation in a reasonable bound, we constrain it within ${\varepsilon}$ by ${{l_{\mathbb{P}}}}$-norm, the function is expressed as follows,
$f({X_{adv}})\neq Y,s.t.{\left\|\delta\right\|_{\mathbb{P}}}<\varepsilon$ (1)
In contrast to traditional methods of generating adversarial samples, adversarial patches are like stickers that are placed over specific local regions of an image. In order to investigate the generation mechanism of adversarial patches, we introduce the concept of the energy model and redefine the process of generating adversarial patches for object detection models from an energy perspective. Generally, in energy-based model, the training set ${S}$ is comprised of tuples ${\{({X^{i}},{Y^{i}});i=1...,N\}}$, which ${{X^{i}}}$ represents the input training example and ${{Y^{i}}}$ denotes its corresponding ground-truth label. Each learning approach can be described as the process of finding parameters ${G\in\left\{{\cal G}\right\}}$ that minimize the energy loss function ${\min L(G,S)}$. However, our adversarial training differs from the training processes of object detection models or energy-based models. The size and distribution of pixel values in the adversarial patch serve as the parameters ${G}$ that require updating. The energy loss function for constructing adversarial patches on the energy-based model is defined by the following equation,
$L_{\text{energy }}(p,S)=\min\mathbb{E}(f(Y,T(p,X)))+R(p)$ (2)
${T(\cdot)}$ represents the transformation function responsible for adding the adversarial patch to the input image, ${p\in{\mathbb{R}^{H\times W\times 3}}}$ denotes the adversarial patch, and ${R(p)={\left\|p\right\|_{\mathbb{P}}}{\text{ = (}}\sum\limits_{i,j}{{{\left|{{p_{(i,j)}}}\right|}^{\mathbb{P}}}}{{\text{)}}^{1/\mathbb{P}}}}$ is the regularization term that restricts the size of each pixel ${{p_{(i,j)}}}$ within the adversarial patch by imposing a ${{l_{\mathbb{P}}}}$-norm constraint. During the adversarial patch training process, we optimize the confidence scores of candidate bounding boxes belonging to the “person” category, which are used as the objective for the optimization algorithm (i.e., the output of ${f(Y,T(p,X))}$). The energy loss function incorporates the confidence scores of all candidate boxes as part of the optimization process. The adversarial patch serves as the parameter to be updated iteratively until the energy of the target category is minimized through the optimization process.
III-B Generating Adversarial Patch
Transformation Function: To facilitate the generalization of the adversarial patch from the digital to the physical world, we leverage the Expectation over Transformation(EoT)[41] framework and introduce a set of physical transformations ${T_{EoT}}$ when applying the patch to an object’s surface. These transformations incorporate variations in contrast, brightness, angles, and other physical environmental factors, thereby ensuring the robustness and effectiveness of the adversarial patch across diverse physical settings. However, in real-world physical environments, the scale of the patch tends to decrease as distance increases and the resulting imagery becomes more blurred. Building on the EoT framework, we introduce a novel distance transformation function. We employ a Gaussian kernel to blur the image, ensuring that the smaller the target object becomes, the larger the convolution kernel becomes, resulting in a more blurred adversarial patch image. Our goal is to develop an adversarial patch pattern that can be printed on a T-shirt to enable a person to evade object detection models. To achieve this goal, we introduce the Thin Plate Splines(TPS)[42] approach to simulate non-rigid deformations such as folds on the T-shirt’s surface. Using the aforementioned transformation methods, we devise a transformation function ${{T_{p}}(\cdot)}$ to apply the adversarial patch to the target image,
$\begin{gathered}{X_{adv}}={T_{p}}(p,X)=\hfill\\ {t_{EoT}}((1-{M_{p}})\cdot X+{t_{d}}({M_{p}}\cdot{t_{TPS}}(p)))\hfill\\ \end{gathered}$ (3)
${{M_{p}}\in{[0,1]^{d}}}$ represents the mask region of the adversarial patch on the target image, which is determined based on the bounding box obtained from the label. The ${{T_{p}}(\cdot)}$ function encompasses three main types of transformations, of which the first one involves applying TPS transformation(${{t_{TPS}}\in{T_{TPS}}}$) to the generated adversarial patch. This serves to introduce distortions, thereby simulating non-rigid deformations on clothing surfaces. Subsequently, based on the size of the bounding box in the input image, the distance is estimated and a distance transformation function(${{t_{d}}\in{T_{d}}}$) is applied to the generated adversarial patch. This enables the patch image to be blurred to varying degrees. In the last step, a regular physical transformation(${{t_{EoT}}\in{T_{EoT}}}$) is applied to the adversarial patch to replicate typical variations that occur in physical environments. The complete process is illustrated in Fig.3,
Figure 3: The transformation process of the adversarial patch.
Adversarial Loss: For contemporary object detection models, it is crucial to design an appropriate energy function that can effectively deactivate the detection of the person category by minimizing the energy. Given images ${X}$ and the ground-truth labels ${Y}$, contemporary object detectors take this information as input, and utilize detection models to generate bounding boxes’ information and corresponding confidence scores for the target category. The candidates with scores exceeding a predefined threshold are selected as the predicted results. Consequently, we incorporate the expected confidence scores of the output as a part of the energy function for adversarial patch optimization, aiming to reduce the target’s confidence scores by optimizing the adversarial patch. Specifically, we begin by randomly initializing the adversarial patch and utilizing the transformation function ${{T_{p}}(\cdot)}$ to apply the patch to the original image, which serves as the input for the object detection models. Previous research incorporated the maximum confidence score of the target category, which was output by the detection network, as a constituent element in the loss function. Unlike previous work[18, 43], we consider all ${N}$ bounding boxes output by the detection model as optimization targets during each iteration, taking into account changes of all confidence scores comprehensively. To reduce the computational cost, we select candidate boxes whose confidence scores for the category “person” are higher than the set threshold value ${\mu}$ as attack objects. Subsequently, we continually optimize and update the patch pixels through gradient descent. In the case of one-stage object detection models, such as YOLOv2, YOLOv3, the detection network directly outputs all bounding boxes and associated category information. We integrate confidence scores of candidate boxes relevant to the “person” category as a constituent of the energy function. For two-stage object detection models like Faster R-CNN, person-related bounding boxes are extracted through Region Proposal Network(RPN), incorporating both positive and negative sample information. Positive candidate boxes with confidence scores surpassing a threshold are selectively chosen as attack targets. Hence, we compute the confidence score of all bounding boxes in the adversarial sample that exceeds the threshold ${\mu}$ of the target detection network. This score is merged into the energy function used for training the adversarial patch. Through iterative optimization, the patch is fine-tuned to minimize the total energy loss, as defined in Eq.4. The generative algorithm, designed based on the energy function, optimizes the adversarial patch during training, thereby reducing energy and causing a malfunction in the target detector’s person detection capability. By training a highly complex adversarial patch, the energy of the remaining categories is elevated, leading to a reduction in the energy associated with the person category. Sec.IV-D of the experiment provides a more detailed explanation regarding the impact of the adversarial patch on the energy levels of various categories.
${L_{obj}}=\mathop{\min\mathbb{E}}\limits_{p}\left(\sum\limits_{j=1}^{N}{\max({f_{j}}(Y,{T_{p}}(p,X)),\mu)}\right)$ (4)
III-C Dynamically Optimized Ensemble Model
To achieve person invisibility across multiple object detection models, the generated adversarial patch effectively impacts the predictive information of all detection algorithms, resulting in a significant reduction of the energy associated with the person category in the output. For ${K}$ object detection models, ${f(\cdot)}$ denotes its outputs and ${\mathbf{w}=[{{\text{w}}_{1}},{{\text{w}}_{2}},...,{{\text{w}}_{K}}]}$ denotes the weights used to adjust different models, subject to constraint ${\sum\limits_{i=1}^{K}{{{\text{w}}_{i}}}=1}$. The ensemble model for the adversarial patch can be expressed as follows,
$\mathop{\min}\limits_{p}\sum\limits_{i=1}^{K}{{{\text{w}}_{i}}\cdot{f_{i}}(p)}$ (5)
However, it is challenging to achieve an ideal balance point for the weights in object detection models with vastly different structures such as YOLO and Faster R-CNN, and overfitting is therefore prone to occur during the training process. Sec.IV-E of this paper provides a detailed analysis of this phenomenon. To address the aforementioned challenges, we propose a dynamically optimized ensemble model for multiple object detection models based on the Min-Max adversarial training approach. During the training process, the proportions of each detection model’s contribution to the overall energy output are dynamically adjusted through parameter tuning to achieve the universality of the adversarial patch, as calculated in Eq.6. The training procedure involves an outer minimization phase, with fixed parameters ${\mathbf{w}}$, where the adversarial patch is optimized through backpropagation to minimize the energy output of the ensemble model. And an inner maximization phase is performed with a fixed adversarial patch, using projected gradient descent to balance the weights ${\mathbf{w}}$ of different models and achieve the maximum energy output of the ensemble model. However, excessive parameter fluctuations can lead to oscillations during the training process, making it difficult to converge to an optimal adversarial patch. On the other hand, if the parameter fluctuations are too small, the loss may stop decreasing after reaching a certain point, failing to reach the desired level. In order to stabilize the process of the weights ${\mathbf{w}}$, the dynamic rate adjustment parameter $\nu$ is increased to control the speed of ${\mathbf{w}}$ change. The value of the parameter $\nu$ was verified in Sec.IV-A.
This method of using a single encoding can compromise the generalization capabilities across other models, which can lead to instability during the training process due to potential training bias towards a specific model [44]. Thus, to alleviate this issue, it is common to introduce a strong concave regularizer [45] in the inner maximization step.
${L_{patch}}=\mathop{\min}\limits_{p}\{\mathop{\max}\limits_{{\mathbf{w}}\in W}\sum\limits_{i=1}^{K}{{{\text{w}}_{i}}L_{obj}^{i}}-L({\mathbf{w}})\}$ (6)
[…]
The training process is illustrated as Algorithm 1. In line 1 we initialize the adversarial patch, parameters, and model. Then, load the dataset and the corresponding labels in line 3. In line 4, we overlay the adversarial patch $p$ onto the original sample $x$ through the transformation function ${{T_{p}}(\cdot)}$ to obtain the adversarial sample ${x_{adv}}$. The 5-6 lines enter the external minimization stage, the fixed weight parameter ${\mathbf{w}}$ is the value in the ($t$-1) state, and the total output energy loss of different target detection models is calculated. In line 7, we update the adversarial patch through backpropagation to minimize the total output energy loss. The 8 line is the internal maximization stage, the fixed parameter $p$ is the value in the ($t$) state, and ${\mathbf{w}}$ is updated by projected gradient descent to maximize the total output energy loss. Finally, repeat the Min-Max confrontation training until the maximum number of iterations is reached.
Algorithm 1 Dynamically optimized ensemble model
0:  clean samples $X$, ground-truth labels $Y$, Physical transformation $T$, $K$ object detection models $F=\{{f^{1}},...,{f^{K}}\}$, bounding boxes output by the detection model $N$.
0:  adversarial patch ${p_{adv}}$
1:  Initializing object detection models $F$, setting the initial weight parameters ${{\mathbf{w}}^{(0)}}$, regularize coefficient $\gamma$, dynamic rate adjustment parameter $\nu$, initializing adversarial patch $p$ with random values.
2:  for the max epochs do
3:     The mini-batch of samples consisting of $n$ samples, $x\in X$, $y\in Y$
4:     adversarial samples ${x_{adv}}={T_{p}}(p,x)$
5:     outer minimization: fixing ${\mathbf{w}}={{\mathbf{w}}^{(t-1)}}$,
6:     calculating $\sum\limits_{i=1}^{K}{{{\text{w}}_{i}}\{}\mathbb{E}(\sum\limits_{j=1}^{N}{\max(f_{j}^{i}(Y,{T_{p}}(p,X)),\mu)}\}$
7:     updating $p$ by minimizing through backpropagation
8:     inner maximization: fixing $p={p^{(t)}}$, updating weight ${{\mathbf{w}}^{(t)}}$ through projected gradient descent, ${{\mathbf{w}}^{(t)}}=pro{j_{W}}({{\mathbf{w}}^{(t-1)}}+\nu{\nabla_{\mathbf{w}}}(\sum\limits_{i=1}^{K}{{{\text{w}}_{i}}L_{obj}^{i}}-R({\mathbf{w}})))$
9:  end for
III-D Overall Optimization Process
We aim to generate universal adversarial patches that can be effective in both the digital and physical world, with a focus on optimizing the patch’s pixel values and smoothness. The adversarial patch pixel value constraint component optimizes the color disparity between adversarial patches and regular printers by evaluating the Non-Printable Score. This definition is reflected mathematically in the following equation,
${L_{NPS}}=\sum\limits_{{i_{patch}}\in p}{\mathop{\min}\limits_{{c_{print}}\in C}\left|{{i_{patch}}-{c_{print}}}\right|}$ (7)
${c_{print}}$ represents a printable color and ${i_{patch}}$ represents a pixel in the adversarial patch $p$. Devices such as printers and screens can only reproduce a subset of the ${[0,1]^{3}}$ RGB color space. Furthermore, to ensure the smoothness of generated adversarial patches, the smoothness adjustment component calculates the minimization of smoothness loss (${L_{smooth}}$) for the adversarial patches.
${L_{smooth}}=\sum\limits_{i,j}{{{({{({p_{i,j}}-{p_{i+1,j}})}^{2}}+{{({p_{i,j}}-{p_{i,j+1}})}^{2}})}^{\frac{1}{2}}}}$ (8)
${p_{i,j}}$ refers to the pixel value at the coordinate $(i,j)$ in the adversarial patch $p$. When adjacent pixel values are close, the smoothness loss ${L_{smooth}}$ is low (i.e., perturbation is smooth). Therefore, by minimizing ${L_{smooth}}$, the smoothness of the perturbed image can be increased, enhancing the feasibility of physical attacks. Finally, in order to generate adversarial patches that are effective in both the physical and digital worlds, we define the total training energy loss ${L_{attack}}$ as follows.
${L_{attack}}={L_{patch}}+\alpha\cdot{L_{NPS}}+\beta\cdot{L_{TV}}$ (9)
$\alpha$ and $\beta$ denote weight coefficients determined empirically. The goal of the optimization process is to minimize the overall energy function ${L_{attack}}$ by jointly minimizing its constituent terms.
IV Experiment
IV-A Implementation Details
To conduct experiments on digital image attacks, we utilized the Inria Person dataset[46] consisting of 614 person images for training and 288 images for testing purposes. To evaluate the attack effectiveness and transferability of adversarial patches, we first measure their impact by computing the Average Precision (AP) values of the images with and without the patches. The larger the drop in AP, the stronger the attack effectiveness. The adversarial patches designed in this study were mainly intended to deceive mainstream one-stage object detection models, namely YOLOv2, YOLOv3, and YOLOv4. As for more complex two-stage models, these patches were found to result in failure to predict when applied to Faster R-CNN detection models with backbones based on VGG16 and ResNet50 architectures. All detection models were trained on the COCO dataset, with the output confidence threshold set to 0.5 and the Non-Maximum Suppression (NMS) threshold to 0.4 during the inference stage. We consider three different sets of object detection models to design ensemble models for the adversarial patches:
  • DOEPatch(YY): Adversarial patch for ensemble model training composed of YOLOv2 and YOLOv3.
  • DOEPatch(YF): Adversarial patch for ensemble model training composed of YOLOv3 and Faster-VGG16.
  • DOEPatch(YYF): Adversarial patch for ensemble model training composed of YOLOv2, YOLOv3 and Faster-VGG16.
And based on the designed Energy model (EmPatch), we retrained adversarial patches:
  • EmPatch(Y1): Adversarial patch for single model YOLOv2.
  • EmPatch(Y2): Adversarial patch for single model YOLOv3.
  • EmPatch(F): Adversarial patch for single model Faster-VGG16.
The entire algorithm is designed in PyTorch and trained on an NVIDIA RTX 3090 24GB GPU cluster. The adversarial patches are set to a size of $3\times 300\times 300$, optimized using the Adam[47] optimizer, and the learning rate for external minimization optimization is set to 0.03. The weight coefficients were determined based on experience, $\alpha=0.01$, $\beta=0.165$, $\mu=0.4$ and the regularization coefficient was set $\gamma=0.1$. The internal dynamic rate adjustment parameter $\nu$ was determined based on experimental results on DOEPatch(YF), as shown in Fig.4. Validation showed that with $\nu=0.78$, the adversarial patch achieved good attack effects on both YOLOv2 and YOLOv3.
Figure 4: The process of AP value changes with the variation of the internal dynamic rate adjustment parameter $\nu$.
IV-B Digital Attack Evaluations
To more fully demonstrate the superiority of the adversarial patch based on dynamically optimized ensemble model design in terms of attack performance and transferability, this study added six comparative experiments to compare with related work that has publicly released adversarial patch results in recent years. Fig.5(a) is an adversarial patch constructed using random noise, Fig.5(b) is a pure color adversarial patch designed as Blue, Fig.5(c) is an adversarial patch called AdvPatch designed by Thys et al.[18] for the YOLOv2 algorithm, Fig.5(d) is an adversarial patch named TCEGA designed by Hu et al.[21] for the YOLOv2 algorithm, Fig.5(e) is an adversarial patch designed based on StyleGAN to attack the YOLOv2 algorithm named NaPatch[48], and Fig.5(f) is an adversarial patch designed for the Faster R-CNN algorithm named UpcPatch[49]. Based on the energy function reconstruction algorithm for adversarial patch generation, this study designed adversarial patches EmPatch(Y1), EmPatch(Y2), and EmPatch(F) respectively for YOLOv2, YOLOv3, and Faster R-CNN, as shown in Fig.5(g), Fig.5(h), and Fig.5(i), respectively. Adversarial patches designed based on the dynamically optimized ensemble model, DOEPatch(YY), DOEPatch(YF), and DOEPatch(YYF), are shown in Fig.5(j), Fig.5k, and Fig.5(l), respectively. Note that AdvPatch and TCEGA were trained on the same dataset and used the same YOLOv2 model and parameters as those in this study, making them highly comparable. However, NaPatch and UpcPatch focused more on the stealthiness of the adversarial patches and exhibit certain differences compared to this study in terms of the dataset used.
All adversarial patches were applied to the images in the Inria test set, and the effectiveness of the attacks was evaluated by calculating the changes in the AP of the person category in the output of different object detection models. Experimental results in the digital world are shown in Tab.I, where “Clean” represents the AP value of the person category in the original test set. The AP values for the test sets with 12 additional adversarial patches were tested using 5 mainstream object detection models. Under the detection of YOLOv2, DOEPatch(YY) reduced the AP value by 67.49%, slightly lower than AdvPatch. However, the attack performance of DOEPatch(YY) is better than other adversarial patches such as TCEGA. Meanwhile, the best attack performance was achieved under the detection of YOLOv3, where DOEPatch(YY) was able to reduce the AP value to 29.20%. Our energy-based adversarial patch EmPatch(Y2) achieved an AP reduction of 54.11% under the detection of YOLOv3, and its attack performance was far superior to the test results of other patches. For the more complex Faster-VGG detection algorithm, DOEPatch(YF) and DOEPatch(YYF) achieved the highest attack effect, reducing AP values by 43.94% and 44.81%, respectively. They also demonstrated strong attack effectiveness under the detection of YOLOv3, reducing AP values by 41.91% and 41.65%, respectively. Due to differences in the dataset, UpcPatch only reduced the AP value by 6.10% in the case of Faster R-CNN.
Figure 5: (a)-(f) show the comparative adversarial patches, while (g)-(i) show the adversarial patches generated by us for YOLOv2, YOLOv3, and Faster R-CNN, respectively. (j)-(l) show the adversarial patches generated based on three sets of ensemble models.
To comprehensively compare the transferability of adversarial patches, we conducted black-box attack tests on YOLOv4 and Faster-Res detection algorithms. Our trained DOEPatch(YY) can reduce the AP value by 50.55% in the detection of YOLOv4, while AdvPatch and TCEGA can only reduce it by 29.14% and 28.21%, respectively. Under Faster-Res, the attack performance was generally poor, but DOEPatch(YY) and DOEPatch(YF) still achieved the best attack results, reducing the AP value by 18.55% and 17.05%, respectively. Overall, the adversarial patches generated based on a dynamically optimized ensemble model exhibited strong attack performance across multiple targeted models and demonstrated favorable transferability by causing significant output interference against unknown black-box models.
TABLE I: The experiment in the digital world measured the attack performance of various adversarial patches against targeted detectors, with changes in AP values used as the metric. The left axis lists all the adversarial patches, and the top axis lists the object detectors targeted for attack, the best case is in bold.
YOLOv2(drop) YOLOv3(drop) Faster-VGG(drop) YOLOv4(drop) Faster-Res(drop)
Clean 80.68%(-) 87.00%(-) 91.04%(-) 94.60%(-) 90.53%(-)
Noise 76.85%(3.83) 85.95%(1.05) 87.32%(3.72) 91.67%(2.93) 83.37%(7.16)
Blue 76.94%(3.74) 85.28%(1.72) 87.68%(3.36) 90.85%(3.75) 85.89%(4.64)
AdvPatch[18] 12.59%(68.09) 70.08%(16.92) 65.23%(25.81) 65.46%(29.14) 74.17%(16.36)
TCEGA[21] 15.46%(65.22) 57.43%(29.57) 68.23%(22.81) 66.39%(28.21) 76.20%(14.33)
NpPatch[48] 54.71%(25.97) 74.88%(12.12) 83.35%(7.69) 87.54%(7.06) 83.11%(7.42)
EmPatch(Y1) 15.02%(65.66) 59.73%(27.27) 61.06%(29.98) 49.49%(45.11) 73.83%(16.70)
EmPatch(Y2) 65.99%(14.69) 32.89%(54.11) 85.70%(5.34) 84.97%(9.63) 85.82%(4.71)
DOEPatch(YY) 13.19%(67.49) 29.20%(57.80) 60.91%(30.13) 44.05%(50.55) 71.98%(18.55)
EmPatch(F) 75.56%(5.12) 80.28%(6.72) 50.97%(40.07) 86.61%(7.45) 78.16%(12.37)
UpcPatch[49] 76.42%(4.26) 84.06%(2.94) 84.94%(6.10) 90.96%(3.64) 85.28%(5.25)
DOEPatch(YF) 59.64%(21.04) 45.09%(41.91) 47.10%(43.94) 74.23%(20.37) 73.48%(17.05)
DOEPatch(YYF) 27.90%(52.78) 45.35%(41.65) 46.23%(44.81) 70.85%(23.75) 74.95%(15.58)
IV-C Physical Attack Evaluations
To validate the attack efficacy of the adversarial patches in the physical world, we printed them on T-shirts to simulate the target detection process in real-world settings. Specifically, the adversarial patches in the digital world were composed of 300$\times$300 pixels, which were enlarged to 30cm×$\times$30cm and printed onto white cotton T-shirts using offset printing techniques for physical-world testing, as illustrated in Fig.6.
Figure 6: Real-world adversarial T-shirts.
To simulate realistic detection scenarios, we recorded two videos of persons’ movements using a c920e camera in indoor and outdoor environments, respectively. One of the persons wore a T-shirt with the adversarial patch, and the attack efficacy of the patch in the physical world is demonstrated in Fig.7. Both YOLOv2 and YOLOv3 were able to detect all persons in all frames without adversarial T-shirts, across different environments. However, using adversarial T-shirts created with DOEPatch(YY), DOEPatch(YF), and DOEPatch(YYF), another person could not be detected in both indoor and outdoor environments. According to the results, the adversarial patches generated were found to be effective in manipulating the recognition capability of object detection models in the physical world. Specifically, the patches were able to enable individuals wearing adversarial T-shirts to evade detection by object detection models, while also exhibiting robust transferability across multiple object detection models. These findings highlight the potential vulnerability of real-world object detection models to adversarial attacks and emphasize the need for further research to enhance the robustness of such models against potential threats.
Figure 7: In the detection results of indoor and outdoor scenes, the adversarial patches printed on the front and back can affect the object detector recognition.
IV-D Interpretability of the Adversarial Patch
Although most works have studied the interpretability of adversarial examples[50], there is no comprehensive explanation of the attack performance for adversarial patches. Consequently, we offer a more comprehensive and insightful explanatory analysis of the phenomenon of how adversarial patches perturb target detection predictions, utilizing the energy function perspective and the visual interpretation tool of Grad-CAM[30]. Adversarial patches generated by the energy function and the dynamically optimized ensemble model exhibit complex and distinct texture features in both color and feature distribution, resembling traditional adversarial patches. Fig.8(a) shows the detection result of DOEPatch(YYF) in YOLOv2, misidentified as the “teddy bear” category. Fig.8(b) depicts the detection result of DOEPatch(YF) in YOLOv3, misidentified as the “apple” category. Fig.8(c) shows the detection result of EmPatch(F) in Faster R-CNN, misidentified as the “potted plant” category. Merely analyzing partial detection results is insufficient for uncovering the underlying mechanisms of adversarial patches. From the perspective of energy function analysis, the training process reduces the energy to impair the object detection capability of the target detector on the “person” category by optimizing the adversarial patch. However, it is challenging to visualize the energy change directly, and we describe the energy change process by comparing the number of selected category candidate boxes after the screening. Fig.8(d) illustrates the changes in the candidate boxes’ output of YOLOv2 upon adding DOEPatch(YYF) to the test samples. Fig.8(e) shows the transformation of the candidate boxes’ output of YOLOv3 after adding DOEPatch(YF) to the test samples, and Fig.8(f) depicts the changes in the candidate boxes’ output of Faster R-CNN after adding EmPatch(F). The blue bars represent the number of candidate boxes output by the detection network after predicting the original samples, while the orange bars represent the output after detecting the adversarial samples. Specifically, the number of candidate boxes representing the “person” category sharply decreases after adding the adversarial patch, while some other categories’ quantities increase to a certain extent. By training adversarial patches with complicated features that enhance the energy of other categories, the energy of the person category is lowered to a certain extent.
Figure 8: (a)-(c) represent the misrecognition results of adversarial patches in object detection. (d)-(f) depict the results of changes in predicted categories.
Merely analyzing from the energy perspective is not comprehensive enough. In order to further analyze the effects of the adversarial attack, we used the pre-trained VGG16 model on ImageNet and combined it with the Grad-CAM algorithm to demonstrate the changes in the attention maps of the target images before and after the attack. Two main changes in the results were observed from the experiments conducted on the test dataset. As shown in Fig.9, the (i) row represents the original sample and the adversarial sample after adding the adversarial patch, the (ii) row represents the recognition results under the detector, and the (iii) row represents the detection results under Grad-CAM. As shown in (ii), in Result 1, the person was not recognized as it was not attended to due to the attention being distracted by the adversarial patch. The feature distribution of the person was altered due to the presence of the adversarial patch. In Result 2, the adversarial patch was misrecognized, and all the attention of the model was attracted to the patch, leading to a change in the model’s decision. Overall, the adversarial patch has the ability to affect the distribution of the target features in the detection network, thereby interfering with the prediction results.
Figure 9: The attention distribution maps of the images before and after our attack. After adding the adversarial patches, the original distribution is disrupted.
IV-E Ablation Studies
To validate the effectiveness of the proposed dynamically optimized ensemble model, we conducted the following two experiments compared to the Average Ensemble model. We designed ensemble models utilizing the one-stage algorithms YOLOv2 and YOLOv3 and obtained respective adversarial patches, DOEPatch(YY) and AEPatch(YY). Additionally, we conducted experiments targeting Faster R-CNN and YOLOv3, which yielded novel adversarial patches, AEPatch(YF) and DOEPatch(YF). The weight parameters, denoted by ${{\text{w}}_{0}}$ and ${{\text{w}}_{1}}$, represent the weights before the detection model, which are constantly optimized during training in the dynamically optimized ensemble model. We manually adjusted the weight parameters of the average ensemble model and conducted five sets of experiments on AEPatch(YF), marked with numerical identifiers, as shown in Tab.II. We evaluated the changes in AP across different detection models, and the detection results are presented in Tab.III.
TABLE II: AEPatch under different settings of ${{\text{w}}_{0}}$ and ${{\text{w}}_{1}}$.
${{\text{w}}_{0}}$ ${{\text{w}}_{1}}$
AEPatch-1(YF) 0.5 0.5
AEPatch-2(YF) 0.2 0.8
AEPatch-3(YF) 0.15 0.85
AEPatch-4(YF) 0.1 0.9
AEPatch-5(YF) 0.05 0.95
TABLE III: The results of the comparative experiments between the two sets.
YOLOv2 YOLOv3 Faster R-CNN
DOEPatch(YY) 13.19% 29.20% 60.91%
AEPatch(YY) 15.62% 36.27% 62.44%
DOEPatch(YF) 59.64% 45.09% 47.10%
AEPatch-1(YF) 78.41% 82.66% 50.29%
AEPatch-2(YF) 76.20% 82.71% 47.83%
AEPatch-3(YF) 78.01% 82.06% 54.83%
AEPatch-4(YF) 72.39% 66.08% 68.85%
AEPatch-5(YF) 69.88% 40.25% 76.45%
In the first comparative experiment, both ensemble models achieved effective attack results, with DOEPatch(YY) exhibiting slightly better detection results than AEPatch(YY). In the second set of experiments, due to the more complex framework structure of the second-order detection algorithm, it was difficult for the average ensemble model to generate effective adversarial patches for both YOLOv3 and Faster R-CNN. Following manual parameter tuning across five sets, the training results revealed that AEPatch-4(YF) exhibited non-convergent behavior while overfitting was observed in the remaining patches. As shown in Fig.10, AEPatch-1(YF) exhibits a feature distribution that closely resembles the adversarial patch, EmPatch(F), trained solely on Faster R-CNN, whereas AEPatch-5(YF) displays a feature distribution more akin to that of the adversarial patch, EmPatch(Y2), trained independently on YOLOv3. Alterations to weight parameters result in shifts in the feature distributions of the adversarial patches. Achieving a balance point between the two distinct feature distributions is challenging through manual parameter tuning. However, DOEPatch(YF), which is generated by dynamic ensembling, addresses this issue and achieves favorable attack performance on both detection models. Therefore, the proposed algorithm addresses common issues of non-convergence and overfitting in ensemble model design, eliminates the complicated process of manual parameter tuning, and generates adversarial patches that can attack multiple detection models simultaneously.
Figure 10: The results of AEPatch(YF) after changes to its parameters.
V Conclusion
In this paper, we propose a universal adversarial patch generation method based on dynamically optimized ensemble models, which can effectively attack multiple mainstream object detection models. Specifically, we design a transformation function for the adversarial patch that simulates the changes in patterns on clothing under real-world conditions. Then, we propose an adversarial patch generation algorithm for object detection models based on energy models. Finally, to enable the transferability of the adversarial patch across multiple detection models, we incorporate the concept of adversarial training and construct a dynamically optimized ensemble model. Through extensive comparative experiments, we demonstrate that our designed adversarial patch exhibits stronger attack performance and better transferability, with high attack effectiveness observed in physical-world scenarios. As the study of large-scale models continues to grow in popularity, it is critical to investigate the potential for a universal adversarial patch that can effectively target all mainstream object detection models. Our proposed algorithm provides new insights for future research on universal adversarial patches.